An electron moving along the x axis has a position
Answer #1

7.- We have that the position of the electron is given by

x=10te^{-3t}

Then the velocity of the electron is given by the derivate of the position function

v=\frac{dx}{dt}

v=\frac{d}{dt}\left(10te^{-3t} \right )......................By the derivate of the product

v=10e^{-3t}+(-3)10te^{-3t}

v=(1-3t)10e^{-3t}

Then we want to know when the velocity is zero

v=0

0=(1-3t)10e^{-3t}......only 1-3t can be zero becasue the e^{-3t} is allways different to zero

0=(1-3t)

3t=1

t=\frac{1}{3}.........................The unit of -3 must be \rm s^{-1}, becaues the argument of e^{()} cannot have unit

t=\frac{1}{3}\rm\;s

Replacinf that time in the equation of the position

x=10\left(\frac{1}{3}\rm\;s \right )e^{-3(1/3\rm \;s)}......The unit of 10 must be \rm \;m/s

x=1.23\rm\;m

8.- The intial velocitu in m/s is

v_o =56.0\rm \; km/h \left(\frac{1000\rm\;m}{1\rm\;km} \right )\left(\frac{1\rm\;h}{3600\rm\;s} \right )

v_o=15.6\rm\;m/s

a) Here we go to ude the equation of the position in a motion with constant acceleration

\Delta x= v_ot+\frac{1}{2}at^2...from this equation

\frac{1}{2}at^2=\Delta x- v_ot

a=\frac{2(\Delta x- v_ot)}{t^2}

a=\frac{2[(20\rm\;m)- (15.6\rm\;m/s )(2.39\; s)]}{(2.39\; \rm s)^2}

a=-6.05\rm\;m/s^2

The acceleration is negative because is deceleration. But its magnitude is

|a|=6.05\rm\;m/s^2

b)Now we have the acceleration of the motion. Then we can find the final velocity is given by the equation

v_f=v_o+at

v_f=15.6\rm\;m/s+(-6.05\rm\;m/s^2)(2.39\rm \;s)

v_f=1.14\rm\;m/s